Hello, I need some help here. The following problem is getting me all messed up and I am not sure what I am doing wrong. The problem is: I have to
Solve the system of equations using the substitution method.
If the answer is a unique solution, present it as an ordered pair: (x,y). If not, specify whether the answer is "no solution" or "infinitely many solutions" and state how you arrived at that conclusion.
-7x + y = 4
2x + 8y = 0
Please help me!!!!!
Thank you in advance.

Answer by RAY100(1637)

(1) -7x +y =4,,,,,,or y = 4+7x
.
(2) 2x +8y =0
.
subst from (1) into (2),, 2x + 8(4+7x) =0
.
2x +32+56x =0
.
58x =-32
.
x=-32/58 = - .552
.
subst in (1),, y = 4 + 7(-.552) = .138
.
check by subst into originals
.
(1) -7(-.552) +(.138) =4,,,, 3.864 +.138 = 4,,,ok
.
(2) 2(-.552) +8 (.138) = -1.104 +1.104 =0,,,,,ok
.
info, if there were " no solutions" it would mean that the lines were " parallel", and never intersect. If you didn"t recognize this from equal slopes, doing the math would result in a statement that is "always not true" , such as 6=7
.
if there were infinite solutions, the lines would be " coincident" , or lay on top of each other. If this was not recognized the math would result in a statement that is " always true", such as 6=6.